# Designing Decision Support Systems Using Counterfactual Prediction Sets

### v2

@@ 1 Introduction
Throughout the years, one of the main focus in the area of machine learning for decision support has been classification tasks. In this setting, the decision support system typically uses a classifier to predict the value of a ground truth label of interest and a human expert uses the predicted value to update their own prediction [4, 27, 8]. Classifiers have become notably accurate in a variety of application domains such as medicine [21], education [52], or criminal justice [15], to name a few. However, their data-driven predictions are not always perfect [36]. As a result, there has been a flurry of work on helping human experts understand when and how to use the predictions provided by these systems to update their own [33, 51, 48, 26]. Unfortunately, it is yet unclear how to guarantee that, by using these systems, experts never decrease the average accuracy of their own predictions [55, 56, 46, 23].
@@
Very recently, Straitouri et al., [44] have argued that an alternative type of decision support systems may provide such a guarantee, by design. Rather than providing a label prediction and letting human experts decide when and how to use the predicted label to update their own prediction, this type of systems provide a set of label predictions, namely a prediction set, and forcefully ask the experts to predict a label value from the prediction set, as shown in Figure 1.111There are many systems used everyday by experts that, under normal operation, limit experts’ level of agency. For example, think of a pilot who is flying a plane. There are automated, adaptive systems that prevent the pilot from taking certain actions based on the monitoring of the environment. Their key argument is that, if the prediction set is constructed using conformal prediction [50, 1], then one can precisely trade-off the probability that the ground truth label is not in the prediction set, which determines how frequently the systems will mislead human experts [^2] Since these systems do not allow experts to predict a label value if it lies outside the prediction set, if the prediction set does not contain the ground truth label, we know that the expert’s prediction will be incorrect. and the size of the prediction set, which determines the difficulty of the classification task the experts need to solve using the system. However, Straitouri et al. are only able to find the (near-)optimal conformal predictor that maximizes average accuracy under the assumption that the experts’ predictions follow a stylized expert model, which they also use for evaluation. In this work, our goal is to lift this assumption and efficiently find the optimal conformal predictor that maximizes the average accuracy achieved by real experts using such a system.
Our contributions. We start by formally characterizing experts’ predictions over prediction sets constructed using a conformal predictor using a structural causal model (SCM) [34]. Building upon this characterization, we identify the following natural counterfactual monotonicity assumption on the experts’ predictions, which may be of independent interest. If an expert succeeds (fails) at predicting the ground truth label from a prediction set and the set contains the ground truth label, the expert would have also succeeded (failed) had the prediction set been smaller (larger) but had still contained the ground truth label. Then, we use this counterfactual monotonicity assumption and the nested structure of the prediction sets provided by conformal prediction to design very efficient bandit algorithms to find the optimal conformal predictor. In particular, we formally show that, in our setting, a variant of the successive elimination algorithm [42], which we refer to as counterfactual successive elimination, achieves an exponential improvement in regret in comparison with vanilla successive elimination.
Finally, we conduct a large-scale user study with $2{,}751$ human subjects who make $194{,}407$ predictions over $19{,}200$ different pairs of natural images and prediction sets. In our study, we experiment both with a strict and a lenient implementation of our decision support systems. Under the strict implementation, experts can only predict a label value from the prediction set whereas, under the lenient implementation, experts are encouraged to predict a label value from the prediction set but have the possibility to predict other label values. Perhaps surprisingly, our results demonstrate that, under the strict implementation, experts achieve higher accuracy. This suggests that, for decision support systems based on prediction sets, limiting experts’ level of agency leads to greater performance than allowing experts to always exercise their own agency. Further, our results also demonstrate that, for the strict implementation, the conformal predictor found by our counterfactual successive elimination algorithm offers greater performance than that found by the algorithm introduced by Straitouri et al., [44].
@@
Further related work. Conformal predictors are just one among many different types of set-valued predictors [11], i.e., predictors that, for each sample, output a set of label values. In our work, we opted for conformal predictors over alternatives such as, e.g., reliable or cautious classifiers [54, 29, 28, 31], because of their provable coverage guarantees, which allow us to control how frequently our decision support systems will mislead human experts. Except for two notable exceptions by Straitouri et al., [44] and Babbar et al., [3], set-valued predictors have not been specifically designed to serve decision support systems. Within these two exceptions, the work by Straitouri et al., [44] is more related to ours. However, it assumes that the experts’ predictions are sampled from a multinomial logit model (MNL), a classical discrete choice model [20]. On the contrary, in our work, we do not make any parametric assumptions about the distribution the experts’ predictions are sampled from. The work by Babbar et al., [3] studies the lenient implementation of our decision support systems, under which the experts appear to achieve lower accuracy in our human subject study, as shown in Figure 3. However, it considers a conformal predictor with a given coverage probability, rather than optimizing across conformal predictors. [^3] Babbar et al., [3] reduce the size of the prediction sets constructed using conformal prediction by deferring some samples to human experts during calibration and testing. However, since such an optimization can be applied both to the strict and the lenient implementation of our system, and we do not find any reason why it would change the conclusions of our human subject study, for simplicity, we decided not to apply it.
@@
Within the vast literature of multi-armed bandits (refer to Slivkins, [42] for a recent review), our work is most closely related to causal bandits [24, 25, 14] and combinatorial multi-armed bandits [9]. In causal bandits, there is a known causal relationship between arms and rewards, similarly as in our work. However, the focus is on using this causal relationship, rather than counterfactual inference, to explore more efficiently and achieve lower regret. In combinatorial multi-armed bandits, one can pull any subset of arms, namely a super arm, at the same time. Then, the goal is to identify the (near-)optimal super arm. While one could view our problem as an instance of combinatorial multi-armed bandits, our goal is not to identify the optimal super arm but the optimal base arm.
@@
In learning under algorithmic triage, a classifier predicts the ground truth label of a given fraction of the samples and leaves the remaining ones to a human expert, as instructed by a triage policy [30, 12, 13, 32]. In contrast, in our work, for each sample, a classifier is used to build a set of label predictions and the human expert needs to predict a label value from the set.
@@ 2 Decision Support Systems based on Prediction Sets
Given a multiclass classification task where, for each sample, a human expert needs to predict a label $y\in\mathcal{Y}=\{1,\ldots,L\}$, with $y\sim P(Y\,|\,X)$, with a feature vector $x\in\mathcal{X}$, with $x\sim P(X)$, the decision support system $\mathcal{C}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$ helps the expert by automatically narrowing down the set of potential label values to a subset of them $\mathcal{C}(x)\subseteq\mathcal{Y}$, which we refer to as a prediction set, using a set-valued predictor [11]. Here, for reasons that will become apparent later, we focus on a strict implementation of the system that, for any $x\in\mathcal{X}$, forcefully asks the expert’s prediction $\hat{y}\in\mathcal{Y}$, with $\hat{y}\sim P(\hat{Y}_{\mathcal{C}}\,|\,X,\mathcal{C}(X))$, to belong to the prediction set $\mathcal{C}(x)$. More formally, this is equivalent to assuming that $P(\hat{Y}_{\mathcal{C}}=\hat{y}\,|\,X=x,\mathcal{C}(x))=0$ for all $\hat{y}\notin\mathcal{C}(x)$. Refer to Figure 1 for an illustration of the decision support system $\mathcal{C}$.
@@
Given a user-specified parameter $\alpha\in[0,1]$, a conformal predictor uses a pre-trained classifier $\hat{f}(x)\in[0,1]^{L}$ and a calibration set $\mathcal{D}_{\text{cal}}=\{(x_{i},y_{i})\}_{i=1}^{m}$, where $(x_{i},y_{i})\sim P(X)P(Y\,|\,X)$, to construct the prediction sets $\mathcal{C}_{\alpha}(X)$ as follows: [^4] The assumption that $\hat{f}(x)\in[0,1]^{L}$ is without loss of generality. Here, the higher the score $\hat{f}_{y}(x)$, the more the classifier believes the ground truth label $Y=y$.
@@
where $s(x_{i},y_{i})=1-\hat{f}_{y_{i}}(x_{i})$ is called the conformal score [^5] In general, the conformal score $s(x,y)$ can be any function of $x$ and $y$ measuring the similarity between samples. Here, we choose $s(x,y)=1-\hat{f}_{y}(x)$ following Sadinle et al., [38]. , $\hat{f}_{y_{i}}(x_{i})$ is the output of the classifier (e.g., the softmax score) for feature vector $x_{i}$ and label value $y_{i}$, and $\hat{q}_{\alpha}$ is the $\frac{\lceil(m+1)(1-\alpha)\rceil}{m}$ empirical quantile of the conformal scores $s(x_{1},y_{1}),\ldots,s(x_{m},y_{m})$. Here, note that, for any sample with feature vector $x$, the prediction sets are nested with respect to the parameter $\alpha$, i.e., $\mathcal{C}_{\alpha}(x)\subseteq\mathcal{C}_{\alpha^{\prime}}(x)$ for any $\alpha>\alpha^{\prime}$. Moreover, the conformal predictor enjoys probably approximately correct (PAC) coverage guarantees, i.e., given tolerance values $\delta,\epsilon\in(0,1)$, we can compute the minimum size $m$ of the calibration set $\mathcal{D}_{\text{cal}}$ such that, with probability $1-\delta$, it holds that [49]
@@
where $(1-\alpha)$ is called the (user-specified) coverage probability. [^6] Most of the literature on conformal prediction focuses on marginal coverage rather than PAC coverage. However, since we will optimize the performance of our system with respect to $\alpha$, we cannot afford marginal coverage guarantees, as discussed in Straitouri et al., [44]. Then, we can conclude that, with high probability, for a fraction $(1-\alpha)$ of the samples, $\mathcal{C}_{\alpha}(x)$ contains the ground truth label and thus cannot mislead the expert—if the expert would succeed at predicting the ground truth label $y$ of a sample with feature vector $x$ on her own, she could still succeed using $\mathcal{C}_{\alpha}$ because $\mathcal{C}_{\alpha}(x)$ contains the ground truth label. On the flip side, for a fraction $\alpha$ of the samples, we know that, if the expert uses $\mathcal{C}_{\alpha}$, she will fail at predicting the ground truth label. [^7] There may be some feature vectors $x$ for which, in principle, a conformal predictor may return an empty prediction set $\mathcal{C}_{\alpha}(x)$. However, during deployment, one can trivially conclude that they will not contain the ground truth label. Hence, in those cases, $\mathcal{C}_{\alpha}$ may allow the expert to choose from $\mathcal{Y}$ (or any other subset of labels). Further, we know that the smaller (larger) the value of $\alpha$, the larger (smaller) the size of $\mathcal{C}_{\alpha}(x)$, and thus the higher (lower) the difficulty of the classification task the expert must solve [53, 6, 7].
The important point above is that, since $\alpha$ is a parameter we choose, we can precisely control the trade-off between how frequently the system misleads the expert and the difficulty of the classification tasks the expert needs to solve. Finally, note that, if we do not forcefully ask the expert to predict a label value from the prediction set, we would not be able to have this level of control and good performance would depend on the expert developing a good sense on when to predict a label from the prediction set.
@@ 3 Prediction Sets through a Causal Lens
In this section, we start by characterizing how human experts make predictions using a decision support system via a structural causal model (SCM) [34], which we denote as $\mathcal{M}$. Our SCM $\mathcal{M}$ is defined by the following assignments:
@@
where $\mathrm{A}$, $U$ and $V$ are independent exogenous random variables and $f_{\mathcal{C}}$, $f_{\hat{Y}}$, $f_{X}$ and $f_{Y}$ are given functions. The exogenous variables $\mathrm{A}$, $U$ and $V$ characterize the (user-specified) coverage probability, the expert’s individual characteristics, and the data generating process for the feature vectors $X$ and ground truth labels $Y$, respectively. The function $f_{\mathcal{C}}$ is a set-function directly defined by the conformal predictor, i.e., $f_{\mathcal{C}}(X=x,A=\alpha)=\mathcal{C}_{\alpha}(x)$, where the calibration set $\mathcal{D}_{\text{cal}}$ is given and thus it does not appear explicitly as an independent variable. Further, as argued elsewhere [34], we can always find a distribution for the exogenous variables $U$ and $V$ and functions $f_{X}$, $f_{Y}$ and $f_{\hat{Y}}$ such that the observational distributions $P(X)$, $P(Y|X)$ and $P(\hat{Y}_{\mathcal{C}_{\alpha}}\,|\,X,\mathcal{C}_{\alpha}(X))$ of interest, defined in the previous section, are given by the distribution $P^{\mathcal{M}}$ entailed by the SCM $\mathcal{M}$, i.e., $P(X=x)=P^{\mathcal{M}}(X=x)$, $P(Y=y\,|\,X=x)=P^{\mathcal{M}}(Y=y\,|\,X=x)$ and
@@
where $\text{do}(\mathcal{C}_{\mathrm{A}}(X)=\mathcal{C}_{\alpha}(x))$ denotes a (hard) intervention in which the first assignment in Eq. 2 is replaced by the value $\mathcal{C}_{\alpha}(x)$. Here, note that we model the stochasticity in both features $X$ and labels $Y$ through the exogenous random variable $V$, instead of considering $X=f_{X}(V)$ and $Y=f_{Y}(X)$, to allow for both causal and anticausal features $X$ [39]. In what follows, we will use the above SCM $\mathcal{M}$ to formally reason about the predictions $\hat{Y}_{\mathcal{C}_{\alpha}}$ made by a human expert under different support systems $\mathcal{C}_{\alpha}$ and then introduce two natural monotonicity assumptions from first principles.
Given a sample with feature vector $x$, we can first conclude that, since the expert’s predictions only depend on the (user-specified) coverage probability $(1-\alpha)$ through the prediction set $\mathcal{C}_{\alpha}(x)$, it must hold that, for any pair of decision support systems $\mathcal{C}_{\alpha}$ and $\mathcal{C}_{\alpha^{\prime}}$ such that $\mathcal{C}_{\alpha}(x)=\mathcal{C}_{\alpha^{\prime}}(x)$, if we observe that the expert has predicted $\hat{Y}_{\mathcal{C}_{\alpha}}=\hat{y}$ using $\mathcal{C}_{\alpha}$, we can be certain that she would have predicted $\hat{Y}_{\mathcal{C}_{\alpha^{\prime}}}=\hat{y}$ had she used $\mathcal{C}_{\alpha^{\prime}}$ while holding “everything else fixed” [34]. Next, motivated by prior empirical studies in the psychology and marketing literature [10, 40, 19, 22], which suggest that increasing the number of alternatives in a decision making task increases its difficulty, we hypothesize and later on empirically verify (refer to Figure 6 in Appendix C) that the following interventional monotonicity assumption holds:
@@
[…]
@@
Moreover, we further hypothesize that the following natural counterfactual monotonicity assumption, which is a sufficient condition for interventional monotonicity, may also hold:

### v3

@@ 1 Introduction
Throughout the years, one of the main focus in the area of machine learning for decision support has been classification tasks. In this setting, the decision support system typically uses a classifier to predict the value of a ground truth label of interest and a human expert uses the predicted value to update their own prediction (Bansal et al.,, 2019; Lubars and Tan,, 2019; Bordt and von Luxburg,, 2020). Classifiers have become notably accurate in a variety of application domains such as medicine (Jiao et al.,, 2020), education (Whitehill et al.,, 2017), or criminal justice (Dressel and Farid,, 2018), to name a few. However, their data-driven predictions are not always perfect (Raghu et al.,, 2019). As a result, there has been a flurry of work on helping human experts understand when and how to use the predictions provided by these systems to update their own (Papenmeier et al.,, 2019; Wang and Yin,, 2021; Vodrahalli et al.,, 2022; Liu et al.,, 2023). Unfortunately, it is yet unclear how to guarantee that, by using these systems, experts never decrease the average accuracy of their own predictions (Yin et al.,, 2019; Zhang et al.,, 2020; Suresh et al.,, 2020; Lai et al.,, 2021).
@@
Very recently, Straitouri et al., (2023) have argued that an alternative type of decision support systems may provide such a guarantee, by design. Rather than providing a label prediction and letting human experts decide when and how to use the predicted label to update their own prediction, this type of systems provide a set of label predictions, namely a prediction set, and forcefully ask the experts to predict a label value from the prediction set, as shown in Figure 1.111There are many systems used everyday by experts that, under normal operation, limit experts’ level of agency. For example, think of a pilot who is flying a plane. There are automated, adaptive systems that prevent the pilot from taking certain actions based on the monitoring of the environment. Their key argument is that, if the prediction set is constructed using conformal prediction (Vovk et al.,, 2005; Angelopoulos and Bates,, 2023), then one can precisely trade-off the probability that the ground truth label is not in the prediction set, which determines how frequently the systems will mislead human experts [^2] Since these systems do not allow experts to predict a label value if it lies outside the prediction set, if the prediction set does not contain the ground truth label, we know that the expert’s prediction will be incorrect., and the size of the prediction set, which determines the difficulty of the classification task the experts need to solve using the system. However, Straitouri et al. are only able to find the (near-)optimal conformal predictor that maximizes average accuracy under the assumption that the experts’ predictions follow a stylized expert model, which they also use for evaluation. In this work, our goal is to lift this assumption and efficiently find the optimal conformal predictor that maximizes the average accuracy achieved by real experts using such a system.
@@
1.1 Our Contributions
We start by formally characterizing experts’ predictions over prediction sets constructed using a conformal predictor using a structural causal model (SCM) (Pearl,, 2009). Building upon this characterization, we identify the following natural counterfactual monotonicity assumption on the experts’ predictions, which may be of independent interest. If an expert succeeds (fails) at predicting the ground truth label from a prediction set and the set contains the ground truth label, the expert would have also succeeded (failed) had the prediction set been smaller (larger) but had still contained the ground truth label. Then, we use this counterfactual monotonicity assumption and the nested structure of the prediction sets provided by conformal prediction to design very efficient bandit algorithms to find the optimal conformal predictor. In particular, we formally show that, in our setting, a variant of the successive elimination algorithm (Slivkins,, 2019), which we refer to as counterfactual successive elimination, achieves an exponential improvement in regret in comparison with vanilla successive elimination.
Finally, we conduct a large-scale user study with $2{,}751$ human subjects who make $194{,}407$ predictions over $19{,}200$ different pairs of natural images and prediction sets. In our study, we experiment both with a strict and a lenient implementation of our decision support systems. Under the strict implementation, experts can only predict a label value from the prediction set whereas, under the lenient implementation, experts are encouraged to predict a label value from the prediction set but have the possibility to predict other label values. Perhaps surprisingly, our results demonstrate that, under the strict implementation, experts achieve higher accuracy. This suggests that, for decision support systems based on prediction sets, limiting experts’ level of agency leads to greater performance than allowing experts to always exercise their own agency. Further, our results also demonstrate that, for the strict implementation, the conformal predictor found by our counterfactual successive elimination algorithm offers greater performance than that found by the algorithm by Straitouri et al., (2023).
An open-source implementation of both the strict and the lenient implementation of our system as well as all the data gathered in our human subject study, which we refer to as ImageNet16H-PS, are available at https://github.com/Networks-Learning/counterfactual-prediction-sets.
1.2 Further Related Work
Our work builds upon further related work on set-valued predictors, multi-armed bandits, and learning under algorithmic triage.
@@ 1.2 Further Related Work
Conformal predictors are just one among many different types of set-valued predictors (Chzhen et al.,, 2021), i.e., predictors that, for each sample, output a set of label values. In our work, we opted for conformal predictors over alternatives such as, e.g., reliable or cautious classifiers (Yang et al.,, 2017; Mortier et al.,, 2021; Ma and Denoeux,, 2021; Nguyen and Hüllermeier,, 2021), because of their provable coverage guarantees, which allow us to control how frequently our decision support systems will mislead human experts. Except for two notable exceptions by Straitouri et al., (2023) and Babbar et al., (2022), set-valued predictors have not been specifically designed to serve decision support systems. Within these two exceptions, the work by Straitouri et al., (2023) is more related to ours. However, it assumes that the experts’ predictions are sampled from a multinomial logit model (MNL), a classical discrete choice model (Heiss,, 2016). On the contrary, in our work, we do not make any parametric assumptions about the distribution the experts’ predictions are sampled from. The work by Babbar et al., (2022) studies the lenient implementation of our decision support systems, under which the experts appear to achieve lower accuracy in our human subject study, as shown in Figure 3. However, it considers a conformal predictor with a given coverage probability, rather than optimizing across conformal predictors. [^3] Babbar et al., (2022) reduce the size of the prediction sets constructed using conformal prediction by deferring some samples to human experts during calibration and testing. However, since such an optimization can be applied both to the strict and the lenient implementation of our system, and we do not find any reason why it would change the conclusions of our human subject study, for simplicity, we decided not to apply it.
@@
Within the vast literature of multi-armed bandits (refer to Slivkins, (2019) for a recent review), our work is most closely related to causal bandits (Lattimore et al.,, 2016; Lee and Bareinboim,, 2018; Kroon et al.,, 2022) and combinatorial multi-armed bandits (Chen et al.,, 2013). In causal bandits, there is a known causal relationship between arms and rewards, similarly as in our work. However, the focus is on using this causal relationship, rather than counterfactual inference, to explore more efficiently and achieve lower regret. In combinatorial multi-armed bandits, one can pull any subset of arms, namely a super arm, at the same time. Then, the goal is to identify the (near-)optimal super arm. While one could view our problem as an instance of combinatorial multi-armed bandits, our goal is not to identify the optimal super arm but the optimal base arm. Further, our work also relates to an extensive line of work using multi-armed bandits in real-world applications, which includes, but is not limited to, efficient data collection for clinical trials (Durand et al.,, 2018), dynamic product pricing (Misra et al.,, 2019; Mueller et al.,, 2019), optimal assortment selection (Agrawal et al.,, 2019), and anomaly detection on networks (Ding et al.,, 2019). The survey by Bouneffouf et al., (2020) includes more examples of real-world applications where multi-armed bandits have been used.
@@
In learning under algorithmic triage, a classifier predicts the ground truth label of a given fraction of the samples and leaves the remaining ones to a human expert, as instructed by a triage policy (Mozannar and Sontag,, 2020; De et al.,, 2020, 2021; Okati et al.,, 2021). In contrast, in our work, for each sample, a classifier is used to build a set of label predictions and the human expert needs to predict a label value from the set.
@@ 2 Decision Support Systems based on Prediction Sets
Given a multiclass classification task where, for each sample, a human expert needs to predict a label $y\in\mathcal{Y}=\{1,\ldots,L\}$, with a feature vector $x\in\mathcal{X}$, with $x,y\sim P(X,Y)$, the decision support system $\mathcal{C}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$ helps the expert by automatically narrowing down the set of potential label values to a subset of them $\mathcal{C}(x)\subseteq\mathcal{Y}$, which we refer to as a prediction set, using a set-valued predictor (Chzhen et al.,, 2021). Here, for reasons that will become apparent later, we focus on a strict implementation of the system that, for any $x\in\mathcal{X}$, forcefully asks the expert’s prediction $\hat{y}\in\mathcal{Y}$, with $\hat{y}\sim P(\hat{Y}_{\mathcal{C}}\,|\,X,\mathcal{C}(X))$, to belong to the prediction set $\mathcal{C}(x)$. More formally, this is equivalent to assuming that $P(\hat{Y}_{\mathcal{C}}=\hat{y}\,|\,X=x,\mathcal{C}(x))=0$ for all $\hat{y}\notin\mathcal{C}(x)$. Refer to Figure 1 for an illustration of the decision support system $\mathcal{C}$.
@@
Given a user-specified parameter $\alpha\in[0,1]$, a conformal predictor uses a pre-trained classifier $\hat{f}(x)\in[0,1]^{L}$ and a calibration set $\mathcal{D}_{\text{cal}}=\{(x_{i},y_{i})\}_{i=1}^{m}$, where $(x_{i},y_{i})\sim P(X,Y)$, to construct the prediction sets $\mathcal{C}_{\alpha}(X)$ as follows: [^4] The assumption that $\hat{f}(x)\in[0,1]^{L}$ is without loss of generality. Here, the higher the score $\hat{f}_{y}(x)$, the more the classifier believes the ground truth label $Y=y$.
@@
where $s(x_{i},y_{i})=1-\hat{f}_{y_{i}}(x_{i})$ is called the conformal score [^5] In general, the conformal score $s(x,y)$ can be any function of $x$ and $y$ measuring the similarity between samples. Here, we choose $s(x,y)=1-\hat{f}_{y}(x)$ following Sadinle et al., (2019)., $\hat{f}_{y_{i}}(x_{i})$ is the output of the classifier (e.g., the softmax score) for feature vector $x_{i}$ and label value $y_{i}$, and $\hat{q}_{\alpha}$ is the $\frac{\lceil(m+1)(1-\alpha)\rceil}{m}$ empirical quantile of the conformal scores $s(x_{1},y_{1}),\ldots,s(x_{m},y_{m})$. Here, note that, for any sample with feature vector $x$, the prediction sets are nested with respect to the parameter $\alpha$, i.e., $\mathcal{C}_{\alpha}(x)\subseteq\mathcal{C}_{\alpha^{\prime}}(x)$ for any $\alpha>\alpha^{\prime}$. Moreover, the conformal predictor enjoys probably approximately correct (PAC) coverage guarantees, i.e., given tolerance values $\delta,\epsilon\in(0,1)$, we can compute the minimum size $m$ of the calibration set $\mathcal{D}_{\text{cal}}$ such that, with probability $1-\delta$, it holds that (Vovk,, 2012)
@@
where $(1-\alpha)$ is called the (user-specified) coverage probability. [^6] Most of the literature on conformal prediction focuses on marginal coverage rather than PAC coverage. However, since we will optimize the performance of our system with respect to $\alpha$, we cannot afford marginal coverage guarantees, as discussed in Straitouri et al., (2023). Then, we can conclude that, with high probability, for a fraction $(1-\alpha)$ of the samples, $\mathcal{C}_{\alpha}(x)$ contains the ground truth label and thus cannot mislead the expert—if the expert would succeed at predicting the ground truth label $y$ of a sample with feature vector $x$ on her own, she could still succeed using $\mathcal{C}_{\alpha}$ because $\mathcal{C}_{\alpha}(x)$ contains the ground truth label. On the flip side, for a fraction $\alpha$ of the samples, we know that, if the expert uses $\mathcal{C}_{\alpha}$, she will fail at predicting the ground truth label. [^7] There may be some feature vectors $x$ for which, in principle, a conformal predictor may return an empty prediction set $\mathcal{C}_{\alpha}(x)$. However, during deployment, one can trivially conclude that they will not contain the ground truth label. Hence, in those cases, $\mathcal{C}_{\alpha}$ may allow the expert to choose from $\mathcal{Y}$ (or any other subset of labels). Further, we know that the smaller (larger) the value of $\alpha$, the larger (smaller) the size of $\mathcal{C}_{\alpha}(x)$, and thus the higher (lower) the difficulty of the classification task the expert must solve (Wright and Barbour,, 1977; Beach,, 1993; Ben-Akiva and Boccara,, 1995).
The important point above is that, since $\alpha$ is a parameter we choose, we can precisely control the trade-off between how frequently the system misleads the expert and the difficulty of the classification tasks the expert needs to solve. Finally, note that, if we do not forcefully ask the expert to predict a label value from the prediction set, we would not be able to have this level of control and good performance would depend on the expert developing a good sense on when to predict a label from the prediction set and when to predict a label from outside the set.
@@ 3 Prediction Sets through a Causal Lens
In this section, we start by characterizing how human experts make predictions using a decision support system via a structural causal model (SCM) (Pearl,, 2009), which we denote as $\mathcal{M}$. Our SCM $\mathcal{M}$ is defined by the following assignments:
@@
where $\mathrm{A}$, $U$ and $V$ are independent exogenous random variables and $f_{\mathcal{C}}$, $f_{\hat{Y}}$, $f_{X}$ and $f_{Y}$ are given functions. The exogenous variables $\mathrm{A}$, $U$ and $V$ characterize the (user-specified) coverage probability, the expert’s individual characteristics, and the data generating process for the feature vectors $X$ and ground truth labels $Y$, respectively. The function $f_{\mathcal{C}}$ is a set-function directly defined by the conformal predictor, i.e., $f_{\mathcal{C}}(X=x,A=\alpha)=\mathcal{C}_{\alpha}(x)$, where the calibration set $\mathcal{D}_{\text{cal}}$ is given and thus it does not appear explicitly as an independent variable. Further, as argued elsewhere (Pearl,, 2009), we can always find a distribution for the exogenous variables $U$ and $V$ and functions $f_{X}$, $f_{Y}$ and $f_{\hat{Y}}$ such that the observational distributions $P(X,Y)$ and $P(\hat{Y}_{\mathcal{C}_{\alpha}}\,|\,X,\mathcal{C}_{\alpha}(X))$ of interest, defined in the previous section, are given by the distribution $P^{\mathcal{M}}$ entailed by the SCM $\mathcal{M}$, i.e., $P(X=x,Y=y)=P^{\mathcal{M}}(X=x,Y=y)$ and
@@
where $\text{do}(\mathcal{C}_{\mathrm{A}}(X)=\mathcal{C}_{\alpha}(x))$ denotes a (hard) intervention in which the first assignment in Eq. 2 is replaced by the value $\mathcal{C}_{\alpha}(x)$. Here, note that we model the stochasticity in both features $X$ and labels $Y$ through the exogenous random variable $V$, instead of considering $X=f_{X}(V)$ and $Y=f_{Y}(X)$, to allow for both causal and anticausal features $X$ (Schölkopf et al.,, 2012). In what follows, we will use the above SCM $\mathcal{M}$ to formally reason about the predictions $\hat{Y}_{\mathcal{C}_{\alpha}}$ made by a human expert under different support systems $\mathcal{C}_{\alpha}$ and then introduce two natural monotonicity assumptions from first principles.
Given a sample with feature vector $x$, we can first conclude that, since the expert’s predictions only depend on the (user-specified) coverage probability $(1-\alpha)$ through the prediction set $\mathcal{C}_{\alpha}(x)$, it must hold that, for any pair of decision support systems $\mathcal{C}_{\alpha}$ and $\mathcal{C}_{\alpha^{\prime}}$ such that $\mathcal{C}_{\alpha}(x)=\mathcal{C}_{\alpha^{\prime}}(x)$, if we observe that the expert has predicted $\hat{Y}_{\mathcal{C}_{\alpha}}=\hat{y}$ using $\mathcal{C}_{\alpha}$, we can be certain that she would have predicted $\hat{Y}_{\mathcal{C}_{\alpha^{\prime}}}=\hat{y}$ had she used $\mathcal{C}_{\alpha^{\prime}}$ while holding “everything else fixed” (Pearl,, 2009). Next, motivated by prior empirical studies in the psychology and marketing literature (Schwartz,, 2004; Haynes,, 2009; Kuksov and Villas-Boas,, 2010; Chernev et al.,, 2015), which suggest that increasing the number of alternatives in a decision making task increases its difficulty, we hypothesize and later on empirically verify (refer to Figure 6 in Appendix C) that the following interventional monotonicity assumption holds:
@@
$P^{\mathcal{M}\,;\,\text{do}(\mathcal{C}_{\mathrm{A}}(X)=\mathcal{C}_{\alpha}(x))}(\hat{Y}_{\mathcal{C}_{\mathrm{A}}}=Y\,|\,X=x)\geq P^{\mathcal{M}\,;\,\text{do}(\mathcal{C}_{\mathrm{A}}(X)=\mathcal{C}_{\alpha^{\prime}}(x))}(\hat{Y}_{\mathcal{C}_{\mathrm{A}}}=Y\,|\,X=x),$ (3)
@@
Moreover, motivated by the aforementioned empirical studies, we further hypothesize that the following natural counterfactual monotonicity assumption, which is a sufficient condition for interventional monotonicity, may also hold: